### v3

Abstract
We present closed analytical approximations for the pricing of basket options, also applicable to Asian options with discrete averaging under the Black-Scholes model with time-dependent parameters. The formulae are obtained by using a stochastic Taylor expansion around a log-normal proxy model and are found to be highly accurate for Asian options in practice as well as for vanilla options with discrete dividends.
keywords:
Stochastic expansion, Asian option, basket option, Black-Scholes, arithmetic averaging, cash dividends.
1 Introduction
The buyer of a vanilla basket option of strike $K$ and maturity $T$ on underlying assets $(S_{i})_{i=1,...,n}$ receives at maturity $T$ $\max\left(\sum_{i=1}^{n}w_{i}S_{i}(T)-K\right)$ for a call and $\max\left(K-\sum_{i=1}^{n}w_{i}S_{i}(T)\right)$ for a put. The weights $w_{i}$ typically sum to one. The buyer of an Asian call option of strike $K$ and averaging dates $(t_{i})_{i=1,...,n}$ on an asset $S$ receives $\max\left(\sum_{i=1}^{n}w_{i}S(t_{i})-K\right)$. Under the Black-Scholes model with a term-structure of interest rates, dividend yields and volatilities, Asian options with discrete arithmetic averaging can be reduced to an equivalent vanilla basket option with a specific correlation structure. Asian (basket) options are particularly popular among commodity traders, but is also traded regularly in the foreign exchange and in the equity derivatives spaces.
There is no exact closed analytic formula for the pricing of arithmetic Asian or vanilla basket options. The problem has been widely studied. In the context of the Black-Scholes model, early valuation techniques were based on approximating the arithmetic mean by a geometric mean [17], or on moment matching against a lognormal distribution as in Levy [12]. Later, Curran [4] introduced a much more accurate pricing based on geometric conditioning which has been more recently further improved, by exploring better integration cutoffs and conditioning variables [5]. Ju proposed a third-order correction for the Levy moment matching in [9], while Zhou [19] looked at matching four moments with the log extended skew normal distribution. Beyond the Black-Scholes model, but restricted to Asian options, Zhang and Oosterlee propose an efficient method based on the Fourier cosine series for affine models with a known characteristic function in [18].
Here, we consider a Black-Scholes setting with time-dependent parameters, and propose a stochastic expansion using the geometric average as a proxy, following the ideas of Etoré and Gobet, who applied their expansion to price vanilla options with discrete dividends under the piecewise-lognormal model in [6], with further refinements in [11]. We start by deriving the first, second and third-order expansions and then evaluate their accuracy on a sample of previously published cases. This allows for a comparison with other accurate approximations.
2 Reductions to vanilla basket options
2.1 From Asian to basket
Let $F(0,t)$ be the forward price of the underlying asset from valuation time to the date $t$. In the Black-Scholes model with time-dependent parameters we have
$F(0,t)=S(0)e^{\int_{0}^{t}r(s)-q(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}\,,$
with $r,q$ being the instantateneous interest rate and dividend yield and $S(t)$ the underlying asset spot at time $t$. Under the risk-neutral measure $\mathbb{Q}$, the forward price follows
$\mathop{}\!\kern 0.0pt\mathrm{d}F(t,T)=F(t,T)\sigma(t)\mathop{}\!\kern 0.0pt\mathrm{d}W(t)\,,$
where $W$ is a $\mathbb{Q}$-Brownian motion.
The buyer of an arithmetic Asian option of strike $K$ and maturity $T$ on an underlying $S$ with observation dates $(t_{i})_{i=1,...,n}$ with $t_{n}=T$, receives at maturity $\max\left(\sum_{i=1}^{n}w_{i}S(t_{i})-K\right)$ for a call and $\max\left(K-\sum_{i=1}^{n}w_{i}S(t_{i})\right)$ for a put. It may be expressed as a basket option of strike $K$ and maturity $T$ on $n$ underlying assets $(S_{i})_{i=1,...,n}$. The covariance between the assets must match the covariance of $S$ between different observation times, this means [^1] We use the Itô isometry to obtain Equations 1 and 3.
(1) $\rho_{i,j}\sqrt{v_{i}v_{j}}=v_{i\wedge j}\,,$
where $\rho_{i,j}$ denotes the correlation between the assets $S_{i}$ and $S_{j}$, $v_{i}=\int_{0}^{t_{i}}\sigma^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$ being the total variance up to time $t_{i}$ for the asset $S$. Each underlying asset forward to $T$ must match the forwards of the single asset to $t_{i}$, meaning
(2) $F_{i}(0,T)=\mathbb{E}[S_{i}(T)]=\mathbb{E}[S(t_{i})]=F(0,t_{i})=S(0)e^{\int_{0}^{t_{i}}(r(s)-q(s))\mathop{}\!\kern 0.0pt\mathrm{d}s}\,.$
Thus, the arithmetic Asian option on $S$ with observations $t_{i}$ corresponds to a vanilla basket option on $S_{i}$ with forwards to maturity $F_{i}(0,T)$, total variances $v_{i}$ and correlations $\rho_{i,j}$ for $i,j=1,...,n$.
2.2 Asian baskets
This may be extended to Asian baskets with payoff at maturity
$\max\left(\sum_{i=1}^{n}w_{i}\sum_{j=1}^{m}\mu_{j}S_{j}(t_{i})-K,0\right)\,,$
where $(\mu_{j})_{j=1...m}$ are basket weights and $S_{j}$ are the different assets with correlation $\rho_{j,l}$. We may define $n\times m$ assets $\hat{S}_{i+n(j-1)}$ with forward $F_{j}(t_{i})$, total variance $\int_{0}^{t_{i}}\sigma_{j}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$, correlation
(3) $\hat{\rho}_{i+n(j-1),k+n(l-1)}=\rho_{j,l}\frac{\int_{0}^{t_{i}\wedge t_{k}}\sigma_{j}(s)\sigma_{l}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}{\sqrt{\int_{0}^{t_{i}}\sigma_{j}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\int_{0}^{t_{k}}\sigma_{l}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}}\,,$
for $(i,k)\in\{1,...,n\}^{2}$, $(j,l)\in\left\{1,...,m\right\}^{2}$. The Asian basket option is equivalent to a vanilla basket option on $\hat{S}_{i+n(j-1)}$ with weights $w_{i}\mu_{j}$.
2.3 Average strike options
A floating strike Asian (call) option paying at maturity $T$, $\max\left(S(T)-k\sum_{i=1}^{n}S(t_{i})\right)$, with $k$ being a strike percentage, can be expressed as an equivalent fixed strike Asian (put) option [15]. Through this symmetry, the formulae given in this paper may be used to price floating (also known as average) strike options.
3 Expansions
3.1 The proxy
An early technique, first explored by Vorst [17] to estimate the price of Asian options with arithmetic averaging is to approximate it by the geometric average option equivalent, because the geometric average of geometric Brownian motions is another geometric Brownian motion and can thus be priced with the usual Black-Scholes formula. This was extended to the more general cases of vanilla basket options by Gentle [7].
Under the risk-neutral measure $\mathbb{Q}$, the forward price for each asset follows
$\mathop{}\!\kern 0.0pt\mathrm{d}F_{i}(t,T)=F_{i}(t,T)\sigma_{i}(t)\mathop{}\!\kern 0.0pt\mathrm{d}W_{i}(t)\,,$
where $W_{i}$ is a $\mathbb{Q}$-Brownian motion, and we have $\mathop{}\!\kern 0.0pt\mathrm{d}W_{i}\mathop{}\!\kern 0.0pt\mathrm{d}W_{j}=\rho_{i,j}\mathop{}\!\kern 0.0pt\mathrm{d}t$. The payoff at maturity of the basket option reads
$V(T)$ $=\left[\eta\left(\sum_{i=1}^{n}w_{i}S_{i}(T)-K\right)\right]^{+}$
$=\left(\sum_{i=1}^{n}w_{i}F_{i}(0,T)\right)\left[\eta\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-{K}^{\star}\right)\right]^{+}=A\left[\eta\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-{K}^{\star}\right)\right]^{+}$
with $\eta=\pm 1$ for a call (respectively a put) option and
$\tilde{a}_{i}=\frac{w_{i}F_{i}(0,T)}{A(T)}\,,\quad A=\sum_{i=1}^{n}w_{i}F_{i}(0,T)\,,\quad{K}^{\star}=\frac{K}{A}\,,$
$S_{i}^{\star}(t)=\frac{S_{i}(t)}{F_{i}(0,t)}=e^{-\frac{1}{2}\int_{0}^{t}\sigma_{i}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s+\int_{0}^{t}\sigma_{i}(s)\mathop{}\!\kern 0.0pt\mathrm{d}W_{i}(s)}\,.$
We approximate the arithmetic average by the geometric average leading to
(4) $V(0)\approx B(0,T)A\mathbb{E}\left[\left(\eta\left(\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-{K}^{\star}\right)\right)^{+}\\ \right]$
with $a_{i}=\tilde{a}_{i}$ as a specific case.
Let
$G(T)=\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}=e^{\sum_{i=1}^{n}-\frac{1}{2}a_{i}\int_{0}^{T}\sigma_{i}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s+\int_{0}^{T}a_{i}\sigma_{i}(s)\mathop{}\!\kern 0.0pt\mathrm{d}W_{i}(s)}$
We have
(5) $m(T)=\mathbb{E}\left[\ln G(T)\right]$ $=-\frac{1}{2}\sum_{i=1}^{n}a_{i}v_{i}\,,$
(6) $\tilde{\nu}^{2}(T)=\mathrm{Var}\left[\ln G(T)\right]$ $=\sum_{i,j=1}^{n}a_{i}a_{j}\rho_{i,j}\int_{0}^{T}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\,.$
with $v_{i}=\int_{0}^{T}\sigma_{i}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$.
Vorst [17], Gentle [7] propose to adjust the strike price of the geometric average option such that the expectation of the term inside the max function is preserved and use $\tilde{K}^{\star}=K^{\star}-(1-\mathbb{E}\left[G(T)\right])$ instead of $K^{\star}$. A different way to preserve the expectation consists in adjusting the spot side instead of the strike side:
(7) $V(0)\approx B(0,T)A\mathbb{E}\left[\left(\eta\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-{K}^{\star}\right)\right)^{+}\right]$
with $\alpha=\frac{1}{\mathbb{E}[G(T)]}$ and $B(0,T)$ is the discount factor to the payment date. The latter is found to be more accurate in Borovykh [1].
Levy [12] matches not only the mean but also the variance of the arithmetic average by a lognormal distribution. We have
(8) $\nu_{A}^{2}=\ln\left(\mathrm{Var}\left[A(T)\right]\right)=\ln\left(\sum_{i,j=1}^{n}a_{i}a_{j}e^{\rho_{i,j}\int_{0}^{T}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}\right)\,..$
We may adjust the $a_{i}$ used in the geometric mean such that the variance of the geometric mean matches exactly the one of the arithmetic mean by using
(9) $a_{i}=\tilde{a}_{i}\frac{\nu_{A}}{\tilde{\nu}}\,.$
Through Equation 6 and the definition of the factor $\alpha$, we obtain the following formula for our Vorst Geometric (VG) and Vorst Levy (VL) approximations, using respectively $a_{i}=\tilde{a}_{i}$ and $a_{i}$ as defined by Equation 9:
(10) $V(0)\approx A\cdot\mathrm{Black}(1,K^{\star},{\nu}^{2},T)$
where $\mathrm{Black}(F,K,v,T)$ denotes the Black-76 formula applied to the forward $F$, strike $K$ and total variance $v$ to maturity $T$ and $\nu=\tilde{\nu}$ for the geometric approximation, $\nu=\nu_{A}$ for the Levy approximation. We have
(11) $\mathrm{Black}(F,K,v,T)=\eta B(0,T)\left[F\Phi(\eta d_{1})-K\Phi(\eta d_{2})\right]$
with $d_{1}=\frac{1}{\sqrt{v}}\left(\ln\frac{F}{K}+\frac{1}{2}v\right)$, $d_{2}=d_{1}-\sqrt{v}$.
We will adopt the Vorst geometric formula with adjusted forward as proxy for our stochastic Taylor expansion. The Vorst geometric formula with strike shift would lead to slightly more complicated equations and exhibit a worse accuracy.
The Vorst Levy formula corresponds exactly to the Levy approximation, but defines a stochastic process based on the individual underlying asset processes. This will allow to define the stochastic expansions based on this alternative proxy as well.
3.2 First order expansion
Let $h(x)=\max(\eta x,0)$, we apply a Taylor expansion of order-1 on $h$:
(12) $\mathbb{E}\left[B(0,T)h\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-K^{\star}\right)\right]=\mathbb{E}\left[B(0,T)h\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-K^{\star}\right)\right]\\ +\mathbb{E}\left[B(0,T)\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}\right)h^{\prime}\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-K^{\star}\right)\right]\\ +\mathrm{Error}_{2}(h)$
where $\mathrm{Error}_{2}(h)\leq_{c}2\left(||\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}||_{3}\right)^{2}$. We may rewrite the second term using a derivative with regards to $K^{\star}$, and interchange derivation and expectation (see Etoré and Gobet [6]) to obtain
(13) $\mathbb{E}\left[\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}\right)h^{\prime}\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-K^{\star}\right)\right]\\ =-\frac{\partial}{\partial K^{\star}}\mathbb{E}\left[\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}\right)h\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-K^{\star}\right)\right]$
Let $G^{\star}=\alpha G$, we have $\mathbb{E}\left[G^{\star}(T)\right]=1$ and $G^{\star}$ may be thus expressed as
[…]
in probability, with $W_{G}$ being a Brownian motion and $\beta^{2}(s)=\frac{\partial\nu^{2}(s)}{\partial s}$. We interpret $\frac{G^{\star}(T)}{G^{\star}(0)}$ as a change of measure on $\mathcal{F}_{T}$. Under the new induced measure $\mathbb{Q}^{G}$, $\tilde{W}_{G}(T)=W_{G}(T)-\int_{0}^{T}\beta(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$ is a Brownian motion. Then $G^{\star}(T)$ under $\mathbb{Q}^{G}$ has the same law as $G^{\star}(T)e^{\nu^{2}(T)}$ under $\mathbb{Q}$. Thus,
(14) $\mathbb{E}\left[G^{\star}(T)h^{\prime}\left(G^{\star}(T)-K^{\star}\right)\right]$ $=-\frac{\partial}{\partial K^{\star}}\mathbb{E}^{G}\left[h\left(\tilde{G}^{\star}(T)e^{\nu^{2}(T)}-K^{\star}\right)\right]\,.$
The right hand side corresponds to the derivative of a vanilla option price under the Black model with forward $e^{\nu^{2}(T)}$, strike $K^{\star}$ and total variance $\nu^{2}(T)$.
In similar fashion we use a change of measure $\frac{S_{i}^{\star}(T)}{S_{i}^{\star}(0)}$ and define
$\tilde{W}_{i}(t)=W_{i}(t)-\int_{0}^{t}\sigma_{i}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$
to obtain
(15) $\mathbb{E}\left[S_{i}^{\star}(T)h^{\prime}\left(G^{\star}(T)-K^{\star}\right)\right]$ $=-\frac{\partial}{\partial K^{\star}}\mathbb{E}^{i}\left[h\left(\tilde{G}^{\star}(T)e^{\sum_{j=1}^{n}a_{j}\int_{0}^{T}\rho_{i,j}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}-K^{\star}\right)\right]\,.$
We use Equation 14 and Equation 15 in Equation 13 to obtain the first order expansion
(16) $V_{\textsf{VG1}}=A\cdot\mathrm{Black}(1,K^{\star},\nu^{2},T)\\ +A\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\nu^{2}},K^{\star},\nu^{2},T\right)-A\sum_{i=1}^{n}\tilde{a}_{i}\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\bar{v}_{i}},K^{\star},\nu^{2},T\right)\,.$
with
(17) $\bar{v}_{i}=\sum_{l=1}^{n}a_{l}v_{i,l}\,,$ $\quad v_{i,l}=\rho_{i,l}\int_{0}^{T}\sigma_{i}(s)\sigma_{l}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\,.$
In particular, for an Asian, we have $v_{i,l}=\int_{0}^{t_{i}\wedge t_{j}}\sigma^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$. For baskets, the volatilities are taken to be constant equal to each underlying implied volatility to maturity in practice and we have $v_{i,l}=\rho_{i,l}\sqrt{v_{i}v_{l}}$.
The relevant derivatives of the Black-76 formula are given in Appendix A. The formula with the Vorst Levy proxy is exactly the same, but uses different $\nu$ and $a_{i}$.
3.3 Second order expansion
The second-order Taylor expansion involves the additional term
$\frac{1}{2}\frac{\partial}{\partial K^{\star}}\mathbb{E}\left[\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}\right)^{2}h\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-K^{\star}\right)\right]$
We decompose it into three terms. The first term involves only $G^{\star}$, we perform two successive changes of measure around $\frac{G^{\star}(T)}{G^{\star}(0)}$ to obtain
$\mathbb{E}\left[G^{\star}(T)^{2}h(G^{\star}(T)-K)\right]$ $=\mathbb{E}^{G}\left[\tilde{G}^{\star}e^{\nu^{2}(T)}h\left(\tilde{G}^{\star}e^{\nu^{2}(T)}-K^{\star}\right)\right]$
$=e^{\nu^{2}(T)}\mathbb{E}^{G^{2}}\left[h\left(\tilde{\tilde{G}}^{\star}e^{2\nu^{2}(T)}-K^{\star}\right)\right]\,.$
The second term involves the cross product $S_{i}^{\star}(T)G^{\star}(T)$. We perform two successive changes of measure around $\frac{G^{\star}(T)}{G^{\star}(0)}$ and $\frac{S_{i}^{\star}(T)}{S_{i}^{\star}(0)}$ to obtain
$\mathbb{E}\left[S_{i}^{\star}(T)G^{\star}(T)h(G^{\star}(T)-K)\right]=\\ \mathbb{E}^{i}\left[\tilde{G}^{\star}e^{\sum_{j=1}^{n}a_{j}\int_{0}^{T}\rho_{i,j}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}h\left(\tilde{G}^{\star}e^{\sum_{j=1}^{n}a_{j}\int_{0}^{T}\rho_{i,j}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}-K^{\star}\right)\right]=\\ e^{\sum_{j=1}^{n}a_{j}\int_{0}^{T}\rho_{i,j}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}\mathbb{E}^{i,G}\left[h\left(\tilde{\tilde{G}}^{\star}e^{\nu^{2}(T)+\sum_{j=1}^{n}a_{j}\int_{0}^{T}\rho_{i,j}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}-K^{\star}\right)\right]\,.$
The third term involves $S_{i}^{\star}(T)S_{j}^{\star}(T)$. We perform two successive changes of measure around $\frac{S_{i}^{\star}(T)}{S_{i}^{\star}(0)}$ and $\frac{S_{j}^{\star}(T)}{S_{j}^{\star}(0)}$ to obtain
$\mathbb{E}\left[S_{i}^{\star}(T)S_{j}^{\star}(T)h(G^{\star}(T)-K)\right]\\ =\mathbb{E}^{i}\left[\tilde{S}_{j}^{\star}(T)e^{\int_{0}^{T}\rho_{i,j}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}h\left(\tilde{G}^{\star}e^{\sum_{l=1}^{n}a_{l}\int_{0}^{T}\rho_{i,l}\sigma_{i}(s)\sigma_{l}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}-K^{\star}\right)\right]\\ =e^{\int_{0}^{T}\rho_{i,j}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}\mathbb{E}^{i,j}\left[h\left(\tilde{\tilde{G}}^{\star}e^{\sum_{l=1}^{n}a_{l}\left(\int_{0}^{T}\rho_{i,l}\sigma_{i}(s)\sigma_{l}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s+\int_{0}^{T}\rho_{j,l}\sigma_{j}(s)\sigma_{l}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\right)}-K^{\star}\right)\right]\,.$
Our formula for the second order expansion thus reads
(18) $V_{\textsf{VG2}}=A\cdot\mathrm{Black}(1,K^{\star},\nu^{2},T)\\ +A\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\nu^{2}},K^{\star},\nu^{2},T\right)-A\sum_{i=1}^{n}\tilde{a}_{i}\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\bar{v}_{i}},K^{\star},\nu^{2},T\right)\\ +\frac{A}{2}e^{\nu^{2}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{2\nu^{2}},K^{\star},\nu^{2},T\right)+\frac{A}{2}\sum_{i=1}^{n}\tilde{a}_{i}e^{\bar{v}_{i}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\nu^{2}+\bar{v}_{i}},K^{\star},\nu^{2},T\right)\\ -A\sum_{i,j=1}^{n}\tilde{a}_{i}\tilde{a}_{j}e^{v_{i,j}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}},K^{\star},\nu^{2},T\right)\,.$
3.4 Third order expansion
The third order expansion adds the term
$-\frac{1}{6}\frac{\partial}{\partial K^{\star}}\mathbb{E}\left[\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}\right)^{3}h\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-K^{\star}\right)\right]$
In similar fashion as for the second order, we perform three consecutive changes of measure involving $\frac{G^{\star}(0)}{G^{\star}(0)}$, $\frac{S_{i}^{\star}(T)}{S^{\star}(0)}$. This leads to our third order expansion Our formula for the second order expansion thus reads
(19) $V_{\textsf{VG3}}$ $=A\cdot\mathrm{Black}(1,K^{\star},\nu^{2},T)$ $+A\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\nu^{2}},K^{\star},\nu^{2},T\right)-A\sum_{i=1}^{n}\tilde{a}_{i}\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\bar{v}_{i}},K^{\star},\nu^{2},T\right)$ $+\frac{A}{2}e^{\nu^{2}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{2\nu^{2}},K^{\star},\nu^{2},T\right)$ $-A\sum_{i=1}^{n}\tilde{a}_{i}e^{\bar{v}_{i}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\nu^{2}+\bar{v}_{i}},K^{\star},\nu^{2},T\right)$ $+\frac{A}{2}\sum_{i,j=1}^{n}\tilde{a}_{i}\tilde{a}_{j}e^{v_{i,j}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}},K^{\star},\nu^{2},T\right)$ $+\frac{A}{6}e^{2\nu^{2}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{3\nu^{2}},K^{\star},\nu^{2},T\right)$ $-\frac{A}{2}\sum_{i=1}^{n}\tilde{a}_{i}e^{\nu^{2}+2\bar{v}_{i}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{2\nu^{2}+\bar{v}_{i}},K^{\star},\nu^{2},T\right)$ $+\frac{A}{2}\sum_{i,j=1}^{n}\tilde{a}_{i}\tilde{a}_{j}e^{\bar{v}_{i}+\bar{v}_{j}+v_{i,j}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{\nu^{2}+\bar{v}_{i}+v_{j}},K^{\star},\nu^{2},T\right)$ $-\frac{A}{6}\sum_{i,j,l=1}^{n}\tilde{a}_{i}\tilde{a}_{j}\tilde{a}_{l}e^{v_{i,l}+v_{j,l}+v_{i,j}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}+\bar{v}_{l}},K^{\star},\nu^{2},T\right)\,,$
with $\bar{v}_{i}$, $v_{i,j}$ defined in Equation 17.
3.5 Symmetries
We may exploit the symmetries in the double and triple sums to reduce the number terms being summed and speed up the corresponding algorithm. In particular,
$\sum_{i,j=1}^{n}\tilde{a}_{i}\tilde{a}_{j}e^{v_{i,j}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}},K^{\star},\nu^{2},T\right)=\sum_{i=1}^{n}\tilde{a}_{i}^{2}e^{v_{i,i}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{2\bar{v}_{i}},K^{\star},\nu^{2},T\right)\\ +2\sum_{i=1}^{n}\sum_{j=1}^{i-1}\tilde{a}_{i}\tilde{a}_{j}e^{v_{i,j}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}},K^{\star},\nu^{2},T\right)\,.$
and
$\sum_{i,j,l=1}^{n}\tilde{a}_{i}\tilde{a}_{j}\tilde{a}_{l}e^{v_{i,l}+v_{j,l}+v_{i,j}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}+\bar{v}_{l}},K^{\star},\nu^{2},T\right)=\\ 6\sum_{i=1}^{n}\sum_{j=1}^{i-1}\sum_{l=1}^{j-1}\tilde{a}_{i}\tilde{a}_{j}\tilde{a}_{l}e^{v_{i,l}+v_{j,l}+v_{i,j}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}+\bar{v}_{l}},K^{\star},\nu^{2},T\right)\\ +3\sum_{i=1}^{n}\sum_{j=1}^{i-1}\tilde{a}_{i}\tilde{a}_{j}^{2}e^{2v_{i,j}+v_{j,j}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{\bar{v}_{i}+2\bar{v}_{j}},K^{\star},\nu^{2},T\right)\\ +3\sum_{i=1}^{n}\sum_{l=1}^{i-1}\tilde{a}_{i}^{2}\tilde{a}_{l}e^{2v_{i,l}+v_{i,i}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{2\bar{v}_{i}+\bar{v}_{l}},K^{\star},\nu^{2},T\right)\\ +\sum_{i=1}^{n}\tilde{a}_{i}^{3}e^{3v_{i,i}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{3\bar{v}_{i}},K^{\star},\nu^{2},T\right)\,.$
4 Numerical examples for Asian options
4.1 Weekly
We start by reproducing the example of Ju [9, Table 4] of a weekly averaging call option of maturity 3 years with different strikes and volatilities in Table 1.
Table 1: Prices of weekly averaging call options of maturity 3 years with different strikes $K$ and volatilities $\sigma$, with a constant interest rate $r=9\%$.
$(\sigma,K)$ MC Ju LB Deelstra VG1 VG2 VG3 VL3
(0.05, 95) 15.1197 15.1197 15.1197 15.1197 15.1197 15.1197 15.1197 15.1197
(0.05, 100) 11.3069 11.3069 11.3069 11.3069 11.3069 11.3070 11.3069 11.3069
(0.05, 105) 7.5561 7.5562 7.5561 7.5561 7.5561 7.5561 7.5561 7.5561
(0.10, 95) 15.2163 15.2165 15.2161 15.2163 15.2159 15.2163 15.2163 15.2163
(0.10, 100) 11.6390 11.6394 11.6388 11.6390 11.6387 11.6390 11.6390 11.6390
(0.10, 105) 8.3911 8.3913 8.3909 8.3911 8.3908 8.3911 8.3911 8.3911
(0.20, 95) 16.6342 16.6365 16.6322 16.6342 16.6317 16.6341 16.6342 16.6342
(0.20, 100) 13.7626 13.7634 13.7607 13.7626 13.7600 13.7625 13.7626 13.7626
(0.20, 105) 11.2146 11.2134 11.2128 11.2146 11.2118 11.2145 11.2146 11.2146
(0.30, 95) 19.0145 19.0179 19.0089 19.0145 19.0058 19.0140 19.0144 19.0144
(0.30, 100) 16.5766 16.5755 16.5712 16.5766 16.5675 16.5762 16.5766 16.5766
(0.30, 105) 14.3830 14.3774 14.3775 14.3830 14.3733 14.3827 14.3830 14.3830
(0.40, 95) 21.7269 21.7307 21.7148 21.7269 21.7056 21.7256 21.7268 21.7267
(0.40, 100) 19.5738 19.5690 19.5614 19.5738 19.5516 19.5727 19.5737 19.5737
(0.40, 105) 17.6110 17.5978 17.5979 17.6109 17.5878 17.6100 17.6109 17.6108
(0.50, 95) 24.5527 24.5583 24.5296 24.5527 24.5106 24.5498 24.5524 24.5523
(0.50, 100) 22.6115 22.6032 22.5871 22.6116 22.5679 22.6090 22.6113 22.6111
(0.50, 105) 20.8241 20.8023 20.7980 20.8242 20.7791 20.8219 20.8239 20.8238
RMSE 0.0068 0.0115 0.0000 0.0204 0.0011 0.0001 0.0002
MAE 0.0218 0.0262 0.0001 0.0450 0.0029 0.0002 0.0004
It consists in 157 averaging dates, starting at the valuation date $t=0$ and ending at the maturity date $t=T=3$ years. We add the prices for the lower bound and refined conditioned moment matching lognormal approximation of [5] (respectively named ”LB” and ”Deelstra”) as well as the first, second and third order stochastic expansions named ”VLE-1”, ”VLE-2”, ”VLE-3”. The reference prices are obtained by a quasi Monte-Carlo simulation with Brownian-bridge variance reduction on 32 million paths.
The Deelstra lower bound is found to be more accurate than the Curran geometric conditioning in [9]. We also report slightly lower errors for the Ju approximation than what is presented in [9], because the reference prices used in [9] are less precise than ours. The Ju approximation is slightly more accurate than the Deelstra lower bound. The second order stochastic expansion is significantly more accurate, and the third order expansion is of similar accuracy as the more refined Deelstra approximation. The VL proxy leads to slightly larger errors than the VG proxy.
4.2 Monthly
We then consider the example of an Asian option of maturity 1 year with monthly averaging, on the HAL stock ticker, presented in [19]. The interest rate is chosen to be $r=6\%$, the dividend yield $q=0.97\%$, the volatility is $\sigma=41.33\%$ and the spot price is $S=30.78$. The 12 averaging dates are evenly [^2] We reproduce here the example of Zhou as is. It is a manufactured example albeit on realistic market data, as, in practice, the averaging dates are typically not exactly evenly spread out due to holidays and week-ends. spread out. Figure 1 shows the error in price expressed in b.p. of the underlying spot. The reference prices are obtained with a quasi Monte-Carlo simulation on 128 million paths.
Figure 1: Error in the price of Asian options of maturity 1 year on HAL with monthly averaging, for a range of spot moneyness.
The first order expansion is of similar accuracy as the Deelstra lower bound approximation. The second order expansion is already very accurate, with maximum error well below 0.1 b.p., much lower than the error of the Deelstra lower bound approximation. In Zhou [19], their four moments matching is found to be comparable but slightly less accurate than the Deelstra lower bound. The third order and the Deelstra approximations are extremely accurate, below the Monte-Carlo sampling error, as evidenced by the wiggles in the error profile, with however a slightly larger error for the Deelstra approximation. Asian options with weekly averaging lead to similar observations. Again, the VL proxy leads to slightly larger errors than the VG proxy.
4.3 Yearly
Finally, we study the more extreme examples of Lord [13]: yearly averaging call options of maturity 5 and 30 years. The 5 year options are on an asset of spot price $S=100$, with interest rate $r=5\%$, no dividend and and a somewhat large volatility $\sigma=50\%$ while the 30 year option use a volatility $\sigma=25\%$. They respectively consist in 5 and 30 averaging dates. Tables 2 and 3 reproduce [13, Tables 7 and 9] and reports the error in basis points of the underlying spot for three strikes $K$ of Asian moneyness
$M=\frac{K}{\sum w_{i}F(0,t_{i})}-1\,.$
We name LB(2,3) and Deelstra(2,3) the lower bound and conditioned lognormal moment matching approximations in Deelstra et al. [5] using $\delta_{2}$ and $f_{3}$ as defined in their paper. We also test the Deelstra(3,3) variant with $\delta_{3}$ and $f_{3}$.
Table 2: Error in b.p. in the prices of yearly averaging call options of maturity 5 years with different strikes. D($i$,$j$) corresponds to the Deelstra($i$,$j$) approximation with $\delta_{i}$ and $f_{j}$ [5].
Strike $M$ Ju LB(2,3) D(2,3) D(3,3) VG1 VG2 VG3 VL3
58.2370 -0.5 5.99 -7.93 -0.24 -0.43 -10.96 -0.20 0.19 0.31
116.4741 0.0 -2.62 -8.19 -0.03 0.06 -10.67 -0.37 -0.11 -0.14
174.7111 0.5 -9.27 -10.41 -0.21 -0.04 -10.27 0.45 -0.29 -0.23
Table 3: Error in b.p. in the prices of yearly averaging call options of maturity 30 years with different strikes. D($i$,$j$) corresponds to the Deelstra($i$,$j$) approximation with $\delta_{i}$ and $f_{j}$ [5].
Strike $M$ Ju LB(2,3) D(2,3) D(3,3) VG1 VG2 VG3 VL3
118.9819 -0.5 10.30 -3.62 0.05 -0.10 -7.59 -0.51 0.06 0.12
237.9638 0.0 -2.43 -14.04 -0.29 0.05 -7.95 -0.43 -0.04 -0.08
356.9457 0.5 -12.39 -22.87 -0.48 0.03 -7.84 0.02 -0.10 -0.11
For the $(2,3)$ variant, the errors for LB and Deelstra match respectively what is reported in [13] as ”LB_GA” and ”Curran2M+”. The maximum error of the $(3,3)$ variant is slightly larger. The second and third order expansions are found to be nearly as accurate as the Deelstra conditioned lognormal moment matching approximation. The third order here does not appear to improve significantly over the second order. Figure 2 shows that the third-order is markedly more accurate if we widen the moneyness range. Again, the first order expansion exhibits a similar error as the Deelstra lower bound approximation.
Figure 2: Error in the price of Asian options of maturity 5 years with yearly averaging.
For yearly options of maturity 30 years, Table 3 shows the first order expansion to be more accurate than the Ju and the LB(2,3) lower bound approximations. The Deelstra(3,3) performs slightly better than the Deelstra(2,3) on this example. The second order expansion is as accurate as the Deelstra(2,3) approximation within the range of Asian moneyness [-0.5, 0.5] and becomes slightly less accurate for larger Asian moneyness (Figure 3). The third order expansions VG3 or VL3 allow to further reduce the error, but are still less accurate than the best Deelstra(3,3) approximation further away from the money.
Figure 3: Error in the price of Asian options of maturity 30 years with yearly averaging.
4.4 Two remaining observations
We keep the same setting as the maturity 5 years in [13] but change the maturity to 1.1 year such that only two observations remain at $t=0.1$ and $t=T=1.1$. Figure 4 plots the error for a wide range of strikes.
Figure 4: Error in the price of Asian options with observations at $t=0.1$ and $t=1.1$ year.
The Ju approximation leads to a maximum error of 10 b.p., twice as large as the Deelstra lower bound approximations. The first order expansion is slightly less accurate than the Deelstra lower bound and the second order expansion is significantly more accurate (with an maximum error of around 1 b.p.). The third order expansion is nearly as accurate as the Deelstra conditioned moment matching approximations.
5 Numerical examples for basket options
5.1 Accuracy on basket options
Krekel [10] tests various basket approximations in depth. We reproduce and apply some of the more significant tests to our new expansions.
Table 4: Prices of vanilla basket options of maturity $T=5$ years, with homogeneous volatilities $\sigma_{i}=40\%$, varying the correlation $\rho_{i,j}$ for $i\neq j$.
$\rho$ MC Ju MM3 LB Deelstra VG1 VG2 VG3 VL3
0.10 21.692 21.766 21.619 20.124 21.609 20.124 22.224 21.440 21.612
0.30 25.029 25.052 24.994 24.209 24.987 24.209 25.212 24.961 24.985
0.50 28.007 28.013 27.995 27.633 27.990 27.633 28.059 27.994 27.996
0.70 30.743 30.743 30.740 30.620 30.738 30.620 30.752 30.741 30.742
0.80 32.041 32.041 32.041 31.989 32.040 31.989 32.044 32.041 32.041
0.95 33.919 33.919 33.919 33.916 33.919 33.916 33.919 33.919 33.919
RMSE 0.032 0.034 0.741 0.039 0.741 0.231 0.107 0.038
MAE 0.073 0.073 1.568 0.083 1.568 0.532 0.252 0.081
The first test consists in a relatively long term option of maturity 5 years, using the same volatility for each asset and the same cross correlation, but varying this correlation from 10% to 95% (Table 4). We include the results of a three moments matching to the shifted lognormal distribution (named MM3).
This very simple test already puts into evidence that the expansions on the VG proxy are accurate for correlations larger than 50%, the third order being the most accurate, but deteriorate as the correlation decreases to 10%. The expansions on the VL proxy are accurate across the whole range of correlations. Under the settings of this test, the first order expansion is equivalent to the Deelstra lower bound approximation.
Table 5: Prices of vanilla basket options of maturity $T=5$ years, with homogeneous volatilities $\sigma_{i}=40\%$ and correlation $\rho_{i,j}=50\%$, for $i\neq j$, varying the strike $K$.
$K$ MC Ju MM3 LB Deelstra VG1 VG2 VG3 VL3
50.00 54.310 54.310 54.273 54.158 54.300 54.158 54.345 54.290 54.289
60.00 47.481 47.482 47.446 47.270 47.468 47.270 47.524 47.459 47.459
70.00 41.522 41.525 41.491 41.257 41.507 41.257 41.572 41.501 41.502
80.00 36.351 36.355 36.327 36.041 36.335 36.041 36.404 36.332 36.334
90.00 31.876 31.881 31.858 31.530 31.859 31.530 31.930 31.860 31.862
100.00 28.007 28.013 27.995 27.633 27.990 27.633 28.059 27.994 27.996
110.00 24.660 24.667 24.654 24.266 24.644 24.266 24.710 24.651 24.653
120.00 21.762 21.769 21.761 21.356 21.746 21.356 21.808 21.756 21.758
130.00 19.249 19.256 19.252 18.837 19.234 18.837 19.291 19.246 19.248
140.00 17.065 17.072 17.071 16.652 17.052 16.652 17.102 17.065 17.066
150.00 15.164 15.171 15.173 14.753 15.151 14.753 15.196 15.165 15.167
RMSE 0.005 0.021 0.347 0.015 0.347 0.045 0.014 0.013
MAE 0.007 0.036 0.413 0.017 0.413 0.053 0.022 0.021
With a fixed correlation of $50\%$, the third order expansion has the same accuracy as the refined Deelstra approximation across a range of strikes (Table 5). The Ju approximation is the most accurate on this example where the volatilities are, again, homogeneous.
Next we consider at-the-money options, with a constant correlation at 50% but vary the homogeneous volatility from 5% to 100%. In practice, volatilities larger than 50% are exceptional for long term options. This test however puts in evidence the limits of applicability of some of the approximations. The expansions become inaccurate for large volatility $\sigma_{i}>50\%$ for a maturity of 5 years (Table 8). The third order expansion clearly diverges as the volatility is increased. This is true for the two proxies, the VL proxy being more accurate in general for $\sigma_{i}\leq 50\%$ but worse for larger volatilities.
Inhomogeneous volatilities (Table 6) decrease the accuracy of the expansions slightly, but not significantly so, in contrast to the Ju approximation. This explains why the Ju approximation is not so good for Asian options.
Table 6: Prices of vanilla basket options of maturity $T=5$ years, with inhomogeneous volatilities $\sigma_{2}=\sigma_{3}=\sigma_{4}=\sigma$ but $\sigma_{1}=100\%$ and correlation $\rho_{i,j}=50\%$, for $i\neq j$, varying the volatility $\sigma$.
$\sigma$ MC Ju MM3 LB Deelstra VG1 VG2 VG3 VL3
0.05 19.403 35.591 18.513 19.451 19.462 16.579 17.854 18.687 19.251
0.10 21.092 36.189 18.646 20.838 20.998 18.822 19.934 20.542 20.836
0.15 23.001 36.925 18.812 22.602 23.042 21.263 22.286 22.751 22.757
0.20 25.304 37.801 19.016 24.694 25.400 23.836 24.823 25.209 24.987
0.30 30.553 39.967 19.567 29.518 30.572 29.186 30.225 30.541 30.164
0.40 36.023 42.664 20.379 34.722 35.980 34.601 35.841 36.031 35.806
0.50 41.493 45.836 21.608 39.959 41.411 39.920 41.538 41.270 41.283
0.60 46.733 49.393 23.566 45.047 46.746 45.036 47.264 45.719 45.907
0.70 51.961 53.214 26.925 49.880 51.901 49.878 52.998 48.465 48.679
0.80 56.792 57.171 33.186 54.394 56.802 54.394 58.733 47.745 47.711
1.00 65.354 64.932 53.650 62.324 65.627 62.324 70.201 15.447 9.938
RMSE 9.527 15.633 1.610 0.079 2.081 1.726 15.351 16.964
MAE 16.139 25.006 3.086 0.217 3.086 4.791 49.963 55.416
5.2 Asian basket options on realistic market data
Previously, we looked at the accuracy using simple, manufactured market data in order to assess the loss of accuracy when a single parameter was changing. In this section, we consider Asian basket options on the more realistic market data of Beisser [5]. The basket consists in 5 stocks, with different forwards, volatilities and correlations. The averaging is monthly over the last 5 months. Figure 5 shows the error in basis points
$\frac{V_{\textsf{approx}}-V_{\textsf{MC}}}{\sum_{j=1}^{m}\mu_{j}S_{j}(0)}\times 10000\,,$
for three different maturities across a range of forward moneyness
$\frac{K}{\sum_{i=1}^{n}\sum_{j=1^{m}}w_{i}\mu_{j}F_{j}(0,t_{i})}-1\,.$
Figure 5: Error in the price of Asian basket options across a range of strikes, for three different maturities.
Although not displayed, the first order expansion is found to have nearly the same error profile as the Deelstra lower bound approximation. The higher order expansions are accurate for the 6 months and 1 year maturity, but degrade with the 5 years maturity, where the Deelstra conditioned moment matching leads to the smallest error. This is not too surprising in light of the previous tests: the volatility is kept the same for the long term option and for the short term option, so the total variance is significantly larger in the long term case, which degrades the expansions accuracy. If we scale the volatility by the square root of the time to maturity, the error of the 5 years option price becomes very similar to the one of the 1 year option.
6 Numerical examples for vanilla options with cash dividends
Healy [8] shows that the problem of pricing vanilla option under the piecewise constant lognormal model, where the underlying asset jumps from the cash dividend amount at the ex-dividend date, can be reduced to a vanilla basket option pricing problem under the Black-Scholes model. We may thus use our new expansions for basket options to price vanilla options with cash dividends and assess their accuracy against the stochastic expansions of Etoré, Gobet and Le Floc’h [6, 11] and other discrete dividends approximations.
In Table 7, we reproduce the example of a 7-year option, with varying yearly dividend amounts, first presented in [16] and in [6].
Table 7: Price of a call option of maturity 7 years, on an asset with spot price $S=100$ paying yearly dividends of respective amounts 6, 6.5, 7, 7.5 8, 8, 8, starting at $t_{1}=0.9$, with $r=6\%$ and $\sigma=25\%$.
Strike FDM VL3 VL2 LL3 EG3 GS Deelstra
70 27.21395 27.21392 27.21367 27.21769 27.20498 27.26457 27.21395
100 19.48229 19.48226 19.48181 19.48478 19.47843 19.51519 19.48245
130 14.13026 14.13023 14.12969 14.13139 14.12933 14.15233 14.13046
MRE N/A 2E-6 4E-5 1E-4 3E-4 2E-3 1E-5
Figure 6: Error in the price of vanilla options of maturity 10 years with semi-annual dividends for a range of strikes.
The reference, named ”FDM” is the price obtained by the TR-BDF2 finite difference method using 10,000 space steps and 3,650 time-steps and is accurate to the fifth decimal. We also give the values of the third order expansion from Etoré and Gobet [6] named ”EG3”, the third order expansion from Le Floc’h [11] named ”LL3” and the approximation from Gocsei and Sahel [14] named ”GS”.
Figure 6 reproduces the example of a call option of maturity 10 years, on a stock of price $S=100$, paying dividends of amount $2$ on a semi-annual schedule, with the first dividend starting one day from the valuation date at $t_{1}=1/365$ presented in [14]. The interest rate is taken to be $r=3\%$ and the volatility $\sigma=25\%$.
On those two examples, the third order expansion on the Levy proxy is the most precise, and the second order expansion is more accurate than the third order expansions of Etoré, Gobet and Le Floc’h.
7 Conclusion
For Asian options, the second order expansion on the Vorst Geometric proxy is already very accurate, better than the third order formula of Ju or the Deelstra lower bound approximations. The third order expansion is comparable in accuracy to the more sophisticated Deelstra conditioning with moment matching. The stochastic expansions have the advantage of requiring no numerical solver and no numerical integration; they thus allow for a straightforward stable calculation of the greeks, and facilitate a robust implementation. The Vorst Levy proxy leads to a slightly worse accuracy of the prices.
The expansions do not fare as well on basket options. In particular for very large (not so realistic) volatilities, the expansions misbehave, especially the third order expansion. The expansions on the Vorst Geometric proxy are also not so accurate under mild correlations. The Vorst Levy proxy remediates this latter issue. In the context of commodity options, the assets are typically futures on the same commodity but of different maturities, and have a correlation close to one [3]. The payoff also often involves the arithmetic average of those futures prices: the contract is effectively an Asian basket option. Given such a correlation structure, the expansions would behave very well on those options.
When applied to price vanilla European options with cash dividends under the piecewise-lognormal model (by reducing the problem to a vanilla basket option pricing problem), the expansions are found to be the most precise against previously published approximations tailored to the specific case of cash dividends.
Further research may apply similar expansions to Asian spread options, using a double geometric approximation as proxy (see also [2]). Other possible extensions would be for Asian options under discrete dividends assumptions, Asian options under stochastic volatility models with a reasonably fast pricing of European options such as Schöbel-Zhu.
Appendix A Derivatives of the Black-76 formula towards the strike
The following formulae are used to compute the expansions in practice:
(20) $\frac{\partial}{\partial K}\mathrm{Black}(F,K,v,T)$ $=-\eta B(0,T)\Phi(\eta d_{2})\,,$
(21) $\frac{\partial^{2}}{\partial K^{2}}\mathrm{Black}(F,K,v,T)$ $=\frac{B(0,T)}{K\sqrt{v}}\phi(d_{2})\,,$
(22) $\frac{\partial^{3}}{\partial K^{3}}\mathrm{Black}(F,K,v,T)$ $=\frac{B(0,T)}{K^{2}\sqrt{v}}\left(\frac{\phi(d_{2})}{\sqrt{v}}-1\right)\,,$
where $\Phi$, $\phi$ are the cumulative normal distribution function and the normal probability density function.
Appendix B Additional tables
Table 8: Prices of vanilla basket options of maturity $T=5$ years, with homogeneous volatilities $\sigma_{i}=\sigma\%$ and correlation $\rho_{i,j}=50\%$, for $i\neq j$, varying the volatility $\sigma$.
$\sigma$ MC Ju MM3 LB Deelstra VG1 VG2 VG3 VL3
0.05 3.526 3.526 3.526 3.525 3.526 3.525 3.526 3.526 3.526
0.10 7.050 7.050 7.050 7.043 7.050 7.043 7.050 7.050 7.050
0.15 10.570 10.570 10.570 10.548 10.569 10.548 10.570 10.570 10.570
0.20 14.083 14.083 14.083 14.032 14.082 14.032 14.085 14.083 14.083
0.30 21.078 21.080 21.079 20.912 21.072 20.912 21.091 21.078 21.078
0.40 28.009 28.013 27.995 27.633 27.990 27.633 28.059 27.994 27.996
0.50 34.826 34.843 34.730 34.147 34.796 34.147 34.986 34.737 34.750
0.60 41.488 41.519 41.044 40.412 41.450 40.412 41.881 41.070 41.119
0.70 47.940 47.967 46.471 46.390 47.907 46.390 48.768 46.363 46.502
0.80 54.128 54.095 50.433 52.050 54.120 52.050 55.705 48.888 49.139
1.00 65.354 64.932 53.650 62.324 65.627 62.324 70.201 15.447 9.938
RMSE 0.145 3.745 1.281 0.069 1.281 1.547 15.155 16.782
MAE 0.478 11.760 3.086 0.217 3.086 4.791 49.963 55.416